Fuel weighing 230 000 N is pumped into an aircraft.

This fuel moves upwards through a vertical height of 4.7 m.

The power developed by the pump is 1600 W.

Calculate the time needed to refuel the aircraft.


1
Expert's answer
2021-03-03T11:25:22-0500

Calculate the energy needed to lift all fuel:


"E=Fh,"


where F - weight.

Power describes how quickly work can be done (work here is the difference between the final and initial energy of fuel):


"P=\frac{W}{t}=\frac{E-0}{t}=\frac Et,\\\space\\
t=\frac{E}{P}=\frac{Fh}{P}=676\text{ s}."